OpenStudy (anonymous):

which expression is equivalent to (x^2-9)/2x^2 +5x -3

OpenStudy (anonymous):

x^2-9= (x+3)(x-3) 2x^2+5x-3= (x+3)(2x-1) cancel x+3 remains (x-3) over (2x-1)

OpenStudy (anonymous):

(9x^2-12x+4) factor

OpenStudy (anonymous):

(9x^2-12x+4) what are the factors